# CPM: Class-conditional Prompting Machine for Audio-visual Segmentation

### v1

@@ Abstract
Audio-visual Learning Image Segmentation Multi-modal Learning
@@ 4.1.2 Results
We collected the experimental results from the existing benchmark [7] and updated it with recent works [5, 26, 56]. We modified the AVSegFormer [13] with Mask2former [8] and denoted it as AVSegFormer* in the tables to encourage a fair comparison with our method.
Table 1: Quantitative (mIoU, $F_{\beta})$ audio-visual segmentation results (%) for the AVSBench test sets [59, 58] (resized to 224$\times$224) with ResNet50 [16] backbone. Best results in bold, $2^{nd}$ best underlined. Improvements against the $2^{nd}$ best are in the last row.
    D-ResNet50 [16]     Method     AVSBench-Obj.(SS) AVSBench-Obj.(MS) AVSBench-S
    Per-pixel Classification     TPAVI [59]     72.79 84.80 47.88 57.80 20.18 25.20
        AVSBG [15]     74.13 85.40 44.95 56.80 - -
        ECMVAE [38]     76.33 86.50 48.69 60.70 - -
        DiffusionAVS [37]     75.80 86.90 49.77 62.10 - -
        CAVP [7]     85.77 92.86 62.39 73.62 44.70 57.76
    Transformer     CATR [24]     74.80 86.60 52.80 65.30 - -
        AuTR [31]     75.00 85.20 49.40 61.20 - -
        AQFormer [19]     77.00 86.40 55.70 66.90 - -
        AVSegFormer [13]     76.54 84.80 49.53 62.80 24.93 29.30
        AVSegFormer* [13]     79.80 91.41 53.28 64.76 41.64 55.05
        AVSC [29]     77.02 85.24 49.58 61.51 - -
        BAVS [30]     77.96 85.29 50.23 62.37 24.68 29.63
        AVSAC [5]     76.90 86.95 53.95 65.81 25.43 29.71
        QSD [26]     77.60 86.00 59.60 63.50 46.60 -
        COMBO [56]     81.70 90.10 54.50 66.60 33.30 37.30
        CPM     87.64 93.53 65.22 75.22 48.39 64.01
    Improvement     CPM     +1.87 +0.67 +2.83 +1.60 +1.79 +6.25
@@ 4.2 Performance on Low-resolution AVSBench Videos
4.3 Performance on Original AVSBench Videos
4.4 Performance on VPO Dataset
4.5 Ablation Study
4.5.1 Ablation of Key Components
@@ Appendix 0.A Training and Inference Details
Table 6: Quantitative (mIoU, $F_{\beta})$ audio-visual segmentation results (in %) on AVSBench test sets [59, 58] (resized to 224$\times$224) with PVT-V2-B5 [52] backbone. Best results in bold, $2^{nd}$ best underlined. Improvements against the $2^{nd}$ best are in the last row.
    PVT-V2-B5 [52]     Method     AVSBench-Object (SS) AVSBench-Object (MS) AVSBench-Semantics
    Per-pixel Classification     TPAVI [59]     78.74 87.90 54.00 64.50 29.77 35.20
        AVSBG [15]     81.71 90.40 55.10 66.80 - -
        ECMVAE [38]     81.74 90.10 57.84 70.80 - -
        DiffusionAVS [37]     81.38 90.20 58.18 70.90 - -
        CAVP     87.33 93.61 67.31 78.09 48.59 61.97
    Transformer     CATR [24]     81.40 89.60 59.00 70.00 32.80 38.50
        AuTR [31]     80.40 89.10 56.20 67.20 - -
        AQFormer [19]     81.60 89.40 61.10 72.10 - -
        AVSegFormer [13]     82.06 89.90 58.36 69.30 36.66 42.00
        AVSC [29]     80.57 88.19 58.22 65.10 - -
        BAVS [30]     81.96 88.60 58.63 65.49 32.64 36.42
        AVSAC [5]     84.51 91.56 64.15 76.60 36.98 42.39
        QSD [26]     - - - - - -
        COMBO [56]     84.70 91.90 59.20 71.20 42.10 46.10
        CPM     91.26 95.43 68.41 79.09 55.08 69.01
    Improvement     CPM     +3.93 +1.82 +1.10 +1.00 +6.49 +7.04
@@ 0.A.1 Additional Results
0.A.1.1 Results on Resized AVSBench
We follow previous methods [24, 31, 19, 13, 29, 30, 59, 15, 38, 37, 7] to evaluate our model with PVT-V2-B5 [52] backbone on AVSBench-Objects (SS & MS) [59] and AVSBench-Semantics [58] with resized image resolution (224$\times$224). The results in Tab. 6 show that we improve mIoU by $3.93\%$, $1.10\%$ and $6.49\%$ on the respective benchmarks.
@@ 0.A.1.1 Results on Resized AVSBench
0.A.1.2 Average Pooling Vs. Max Pooling
@@
0.A.1.3 Hyper-parameter Analysis
@@ Appendix 0.C Audio Conditional Prompting Visualisation
Appendix 0.D Video Visualisation
We present a qualitative comparison visualization among TPAVI [59], AVSegFormer [13], CAVP [7], and our CPM on AVSBench-Semantics in Fig. 8. Our findings demonstrate that our method offers a more accurate approximation of the true segmentation of objects in the scene compared to alternative methods. For the demonstration of full video examples on AVSBench-Semantics, please refer to the “video_demo.mp4” file within the attached supplementary materials.

### v3

@@ Abstract
Audio-visual Learning Segmentation Multi-modal Learning
@@ 4.1.2 Results
We collected the experimental results from the existing benchmark [7] and updated it with recent works [5, 26, 55]. We modified the AVSegFormer [13] with Mask2former [8] and denoted it as AVSegFormer* in the tables to encourage a fair comparison with our method. Please note that Tab. 1 includes two evaluation protocols. We employ standard semantic segmentation protocols to compute both mIoU and $F_{\beta}$ same as PascalVOC [11], initially mentioned in [58].
Table 1: Quantitative ($\mathcal{J}$, $\mathcal{F}$) audio-visual segmentation results (%) for the AVSBench test sets [58, 57] (resized to 224$\times$224) with ResNet50 [16] backbone. Best results in bold, $2^{nd}$ best underlined. Improvements against the $2^{nd}$ best are in the last row.
    D-ResNet50 [16]     Method     External Model     AVSBench-Semantics (SS) AVSBench-Semantics (MS) AVSBench-Semantics
                $\mathcal{J}$$&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow $\mathcal{J}$$&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow $\mathcal{J}$$&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow
    Per-pixel Classification     TPAVI [58]         78.80 72.79 84.80 52.84 47.88 57.80 22.69 20.18 25.20
        AVSBG [15]         79.77 74.13 85.40 50.88 44.95 56.80 - - -
        ECMVAE [38]         81.42 76.33 86.50 54.70 48.69 60.70 - - -
        DiffusionAVS [37]         81.35 75.80 86.90 55.94 49.77 62.10 - - -
        CAVP         83.84 78.78 88.89 61.48 55.82 67.14 32.83 30.37 35.29
    Transformer     CATR [24]         80.70 74.80 86.60 59.05 52.80 65.30 - - -
        AuTR [31]         80.10 75.00 85.20 55.30 49.40 61.20 - - -
        AQFormer [19]         81.70 77.00 86.40 61.30 55.70 66.90 - - -
        AVSegFormer [13]         80.67 76.54 84.80 56.17 49.53 62.80 27.12 24.93 29.30
        AVSC [29]         81.13 77.02 85.24 55.55 49.58 61.51 - - -
        BAVS [30]         81.63 77.96 85.29 56.30 50.23 62.37 27.16 24.68 29.63
        AVSAC [5]         81.93 76.90 86.95 59.88 53.95 65.81 27.57 25.43 29.71
        QSD [26]         81.80 77.60 86.00 61.55 59.60 63.50 - - -
        COMBO [55]         85.90 81.70 90.10 60.55 54.50 66.60 35.30 33.30 37.30
        CPM         85.92 81.37 90.47 65.40 59.80 71.00 37.05 34.53 39.57
@@ 4.2 Performance on Low-resolution AVSBench Videos
4.3 Performance on Original AVSBench Videos
4.4 Performance on VPO Dataset
@@ 0.A.1 Additional Results
0.A.1.1 Average Pooling Vs. Max Pooling
@@
        Per-pixel Classification     Transformer
   Methods     TPAVI [58]     CAVP [7]     AVSegFormer [13]     CPM
   Training     48 h     22 h     23 h     30 h
   Inference     9 fps     20 fps     7 fps     14 fps
Table 6: Efficiency comparison for training and inference on AVSBench-Semantic [31] using a ResNet50 backbone at original resolution.
0.A.2 Evaluation of Training and Inference Efficiency
We measure the overall training time and frame per second (FPS) required for inference in Tab. 6. The experiments are conducted on the AVSBench-Semantic (AVSS) dataset [57] using a ResNet50 backbone at the original resolution. During inference, we evaluate models on 10 videos, sampling 100 frames per video (i.e., a total of $10\times 100$ frames) to compute the number of frames per second (FPS). The results show that our CPM requires additional training time (e.g., $+8$ hours) and has a slower inference speed (i.e., $-6$ fps) compared to CAVP. However, CPM achieves a $+7$ fps faster inference speed than AVSegFormer for a similar model architecture.
0.A.2.1 Hyper-parameter Analysis